Grade 11Engineering Entrance Exams

A thick rubber rope of length L, density r and Young’s modulus Y is hung from the ceiling of a room. What is the increase in its length due to its own weight?

Jayant Kumar
12 Years agoGrade 11

1 Answer

Vikas TU
6 Years ago
As the weight of wire acts at centre of gravity.
Therefore only half the length of the wire gets extended
Now Y=F/A.(L/2)/△t=Mg(L/2)/A△l
∴△l=ρL^2g/2Y

question mark
Hello,
 
Can we learn and prepare for MHT-CET online ?
engineering entrance exams

Last Activity: 4 Years ago